# Forecasting Treatment Response with Deep Pharmacokinetic Encoders

### v7

@@ 3.4 Models
To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines. We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model. Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al. (2020); Challu et al. (2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al. (2014)), Long-Short Term Memory (LSTM; Sak et al. (2014)), and Temporal Convolution Network with an MLP Decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)). Implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022) and StatsForecast (Garza et al., 2022) libraries. More information on these models and related prior work that leverages these are provided in Appendix A.2. Details on model training and hyperparameter selection are provided in Appendix A.3
3.5 Baselines
We compare our hybrid global-local architecture with the PK encoder to both local and global baselines:
Global Baseline Models.
Models were trained on all patient data to create a single global model:
  • Baseline. Models are trained with the univariate blood glucose signal to demonstrate the forecasting improvement from using exogenous variables, supporting our PK encoder.
  • Baseline Exogenous (Exog.). Models are trained with the univariate blood glucose signals, static features, and sparse exogenous features (CHO and insulin dose values).
@@ 3.5 Baselines
  • Baseline Exogenous (Exog.) - ST. Models are trained with the univariate blood glucose signal, static features, and sparse exogenous features transformed via the sum total (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
Local Baseline Exogenous Models.
Patient-specific baseline exogenous models were compared against our hybrid global-local PK encoder model.
Local PK Encoder Models.
@@
Patient-specific models using the PK encoder were compared against our hybrid global-local PK encoder model.
3.6 Evaluation
@@
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models. Formulations for these metrics are provided in Appendix A.4. We evaluate model performance across all points within the specified time horizon. We also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
@@ 4 Cohort
Simulated data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. Blood glucose time series plots for 6 subjects are shown Fig. 7 in Appendix A.1. Basal and bolus dose distributions for 12 simuluted patients are shown in Fig. 9 in Appendix A.1.
OhioT1DM
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The data includes basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1. Basal and bolus dose distributions for the 12 patients are shown in Fig. 10 in Appendix A.1.
More dataset information is in Appendix A.1.
@@ 5 Results
5.1 Results on Simulated Data
@@
For the simulated dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous model for forecasts evaluated at all values (t-test p-value: 6.6e-6) as shown in as shown in Table 1. For evaluations across all values, the NHITS-PK model provides 3.5% improvement, on average, and up to 9.2% improvement for individual patients over the baseline NHITS sparse exogenous model. For evaluations across critical values, the NHITS-PK model provides 5.2% improvement, on average, and up to 16.4% improvement for individual patients over the baseline NHITS sparse exogenous model. Additionally, the NHITS-PK model outperforms the NHITS-ST model. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 6 in Appendix C.5.
5.2 Results on Real Data
@@
Our hybrid global-local architecture produces models that significantly outperform local models by 15.8%, on average (t-test p-value: 1.7e-5), and up to 31.3% improvement for individual OhioT1DM patients. We show that the NHITS-PK hybrid global-local model outperforms NHITS-PK models developed for individual OhioT1DM patients in Fig. 12. Moreover, this improvement in forecast accuracy is consistent across all patients as shown in Fig. 5 and in Table 4 in Appendix C.1. Additionally, the NHITS-PK model with the hybrid global-local architecture also improves MAE by 14.8%, on average, for patients over the patient-specific NHITS sparse exogenous baseline models (t-test p-value: 1.7e-6).
For the OhioT1DM dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous baseline model for forecasts evaluated at all values (t-test p-value: 5.0e-3) and critical values (t-test p-value: 3.5e-3) as shown in Table 1. For evaluations across all values, The NHITS-PK model provides 1.3%  on average, and up to 3.5% improvement for individual patients over the baseline NHITS sparse exogenous model. For evaluations across critical values, the NHITS-PK provides 1.6%  on average, and up to 4.9% improvement for individual patients over the baseline NHITS sparse exogenous model. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 7 in Appendix C.5. Fig. 13 in Appendix C.2 shows model forecasts during hyperglycemic events for four randomly selected OhioT1DM patients.
Figure 5: The NHITS-PK hybrid global-local model (blue, solid) provides more accurate predictions than local NHITS-PK models (purple, dashed).
Figure 6: In addition to reducing forecasting error during hyperglycemic events, our NHITS-PK encoder enables improved lead time predictions of hypertensive events (blood glucose $\geq$ 180) for 9 of 12 OhioT1DM patients as measured over the 8 trials.
Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
Simulated Dataset
Model All Values Critical Values
Baseline Exog. Baseline Exog.
ETS 9.450 13.457
MLP 10.589 8.433 14.699 10.600
RNN 13.703 20.190 21.433 35.955
LSTM 13.557 10.321 20.537 13.550
TCN 12.096 10.109 17.709 12.660
NHITS 8.268 7.304 11.114 9.094
NHITS-ST 7.055 8.648
NHITS-PK 7.037 8.492
OhioT1DM Dataset
ETS 9.022 10.028
MLP 11.065 11.120 12.630 12.698
RNN 11.325 11.795 13.627 13.552
LSTM 9.811 9.467 11.252 11.178
TCN 11.525 10.760 13.647 12.678
NHITS 9.060 8.873 10.337 10.137
NHITS-ST 8.958 10.131
NHITS-PK 8.758 9.965
@@ 6 Discussion
Hybrid Global-Local Architecture We demonstrate the efficacy of our novel hybrid global-local model, which produces forecasts with significantly smaller errors than patient-specific models across all patients. Our hybrid architecture not only improves accuracy but also enables global models to learn relevant patient-specific information, outperforming global model baselines.
A hybrid global-local model offers practical benefits beyond improved forecasting accuracy as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. The hybrid model can also be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
@@
PK Encoder As a model-agnostic extension, our novel PK encoder significantly reduces forecasting errors in deep learning models for both simulated and real-world data. These improvements are not only statistically significant, but they have potential for positive real-world clinical impact. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts such events earlier, up to 3.6 minutes when compared to the NHITS baseline model, for individual OhioT1DM patients as shown in Fig. 6. This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations. Improving prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
@@
Our results also highlight the importance of using domain knowledge to incorporate exogenous variables into models. Models with sparse features may underutilize valuable information, and feature-engineered representations not informed by domain knowledge, like the ‘sum total’ approach, can underperform compared to baselines. Moreover, Appendix C.4 shows that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset. As such, dose-dependent models, such as proposed in section 3.3, may be needed for more realistic data simulations and models.
Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing or drug selection. It can also be adapted for other medication administration types, such as oral or intravenous injections.
Limitations. We equate $\sigma$ to our k parameter in the log-normal function and hold $\mu$ constant at 1. This simplification reduces the number of potential concentration curve shapes the model can learn. Tuning both $\sigma$ and $\mu$ could further improve forecasts by enabling more flexible concentration curves.
Future Work. Future work will explore fine-tuning global forecasting models with patient-specific parameters for new patients with limited historical data as well as modeling dose-dependent effects using more dynamic nonlinear PK parameter frameworks. Future work will include applying our approach to additional patient cohorts, such as those with T2DM, to further validate our findings.

### v8

@@ 3.4 Models
To assess forecasting performance across various model architectures, we trained global models with 11 different algorithms as baselines. Additionally, to assess our proposed hybrid global-local architecture, we compare the best performing model with the PK encoder to both purely local and global models.
Statistical and Deep Learning Baselines.
@@
We employed Exponential Smoothing (AutoETS; Hyndman 2008) as our statistical baseline model. Our deep learning baselines include models from various architectural categories, such as Linear models, Multi-Layer Perceptron (MLP)-based, Recurrent Neural Network (RNN)-based, Convolutional Neural Network (CNN)-based, and Transformer-based models. We employ the following deep learning models: DLinear (Zeng et al., 2023), MLP (Rosenblatt, 1958), NHITS (Challu et al., 2022), NBEATSx (Oreshkin et al., 2020; Olivares et al., 2022a), RNN (Elman, 1990; Cho et al., 2014), Long-Short-Term Memory (LSTM; Sak et al. (2014)), Temporal Convolutional Network with an MLP decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)), Informer (Zhou et al., 2021), patch time series Transformer (PatchTST; Nie et al. (2023)), and Temporal Fusion Transformer (TFT; Lim et al. (2021)). More information on these models and related prior work is provided in Appendix A.2. Details on model training and hyperparameter selection are provided in Appendix A.3.
Global Model Baselines.
@@
Global models are trained on all patient data to create a single model. Global model baselines include: (1) univariate models trained with only the blood glucose signal and static features, (2) models trained with univariate blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with the univariate blood glucose signal, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022). Univariate models are included as baselines to highlight forecast improvements from training on historical exogenous variables, motivating advanced transformations for sparse, continuous exogenous features using methods like our PK encoder.
With the exception of TFT, most Transformer-based models, including Informer and PatchTST, are not designed to model covariates. For these two models, we can only report global model results for those trained on univariate blood glucose data. Since our proposed PK encoder is intended to supplement model architectures capable of incorporating exogenous variables alongside the endogenous signal, we focus on evaluating the PK encoder in top-performing models that are capable of modeling covariates, such as TFT, NBEATSx, and NHITS. Beyond modeling covariates, residualized MLP-based models like NBEATSx and NHITS are valuable to demonstrate the utility of our PK encoder as NHITS has achieved SOTA performance, outperforming RNNs and Transformers in long-range forecasting tasks while avoiding the quadratic complexity of self-attention in Transformer-based architectures (Challu et al., 2022; Ansari et al., 2024).
Local Model Baselines.
@@
Patient-specific sparse exogenous models as well as patient-specific exogenous models using the PK encoder were compared against our hybrid global-local PK encoder model.
3.5 Model Training
Models were trained using Adam optimizer (Kingma and Ba, 2017). Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as the loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)=\left\{\begin{array}[]{ll}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{array}\right.$ (8)
Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. Optimal hyperparameters were selected based on a validation set with cross-validation. More information on model training and hyperparameters in provided in Appendix A.3.
3.6 Model Evaluation
@@
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE):
$\mathrm{MAE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\frac{1}{H}\sum_{{\tau}=t+1}^{t+H}|\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau}|\quad\text{and}$ (9)
$\mathrm{RMSE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\sqrt{\frac{1}{H}\sum_{\tau=t+1}^{t+H}(\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau})^{2}}.$ (10)
We compute results across all points within the forecast horizon, and present the average results over 8 trials of individually trained models. We also evaluate the performance of the models exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation. To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-test statistics computed on the patients’ mean MAE results.
@@ 4 Cohort
We use simulated and real-world datasets containing patient insulin, CHO time series, and blood glucose measurements. These open-source datasets ensure reproducibility and are widely used in blood glucose forecasting research, featuring some of the largest patient cohorts (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
Simulated data.
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. Appendix A.1 contains examples of blood glucose time series for 6 individuals, shown in Fig. 10, as well as basal and bolus dose distributions for 12 individuals, shown in Fig. 12.
Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower average MAE than NHITS-PK local models (blue, dashed median).
Figure 6: The NHITS-PK hybrid global-local model (orange, solid median) consistently outperforms the local NHITS-PK model (blue, dashed median).
[…]
Simulated Dataset OhioT1DM Dataset
Model All Values Critical Values All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
AutoETS 9.450 13.457 9.022 10.028
DLinear 7.802 7.989 10.575 10.790 11.514 13.011 12.512 14.580
RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
Informer 8.599 12.074 13.714 14.867
PatchTST 7.284 9.842 9.922 10.999
TFT 6.587 5.826 8.518 7.132 9.402 9.197 10.577 10.643
MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
NBEATSx 9.187 6.886 12.757 8.145 9.868 10.280 11.037 11.359
NHITS 8.268 7.304 11.114 9.094 9.060 8.873 10.337 10.137
TFT-PK 5.743 7.042 9.529 10.461
NBEATSx-PK 6.806 7.998 9.983 10.927
NHITS-PK 7.037 8.492 8.758 9.965
Real-world data.
Twelve de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a continuous glucose monitor (CGM) at a frequency of 5-minutes, basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. The sample count and percentages of missing data for each subject are included in Table 4 in Appendix A.1. Appendix A.1 also contains examples of blood glucose time series plots for 6 subjects, shown in Fig. 11, as well as basal and bolus dose distributions for the 12 patients, shown in Fig. 13.
@@ 5 Results
PK-encoder models outperform their sparse exogenous model counterparts.
@@
For simulated data, the TFT-PK model achieves the lowest MAE, on average, among all models for both evaluations across all models and critical values. The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual subjects. For the OhioT1DM dataset, the NHITS-PK model achieves the lowest MAE, on average, among all models for both evaluations across all values and critical values. The PK models, including, TFT-PK, NBEATSx-PK, and NHITS-PK, outperform their sparse exogenous model counterparts by up to 5.2%, 8.5%, 4.9% on evaluations across critical values for individual patients. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.3 Tables 8 and 9, respectively.
Hybrid global-local models significantly improve forecasts over global models.
For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for subjects for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts. For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for subjects for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts. Additionally, Table 2 shows that TFT-PK and NHITS-PK models achieve significantly lower errors, on average—2.4% (p-value: 3.6e-12) and 2.3% (p-value: 4.1e-5), respectively—for subjects in evaluations across all values, compared to their “Sum Total” model counterparts.
Hybrid global-local models significantly improve forecasts over local models.
@@
For the OhioT1DM dataset, the NHITS-PK hybrid global-local model outperforms local NHITS-PK models developed for individual OhioT1DM patients by 15.8%, on average (p-value: 1.7e-5) as shown in Fig. 5. Furthermore, this improvement in forecast accuracy is consistent across all patients as shown in Fig. 6 and Table 3 (standard deviations are provided in Table 6 in Appendix C.1). Our hybrid global-local NHITS-PK model also significantly outperforms local sparse exogenous baseline models by 14.8%, on average, for individual subjects (p-value: 1.7e-6).
Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. Best result is in bold.
Simulated Dataset OhioT1DM Dataset
TFT-PK TFT-ST NHITS-PK NHITS-ST
All Values 5.743 5.885 8.758 8.958
(0.159) (0.187) (0.080) ((0.087)
Critical Values 7.042 7.233 9.965 10.131
(0.270) (0.282) (0.095) (0.100)
Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset. Best results in bold.
ID w/ Exog. (Sparse) w/ Exog. (PK)
Local Global Local Hybrid
540 12.106 10.119 11.880 10.015
544 7.952 7.188 8.067 6.967
552 8.559 7.400 9.132 7.374
559 10.845 9.013 13.051 8.967
563 9.527 8.731 9.427 8.506
567 12.599 9.640 12.132 9.562
570 8.361 7.523 8.466 7.258
575 10.891 9.712 11.095 9.496
584 11.111 9.846 11.518 9.905
588 10.172 8.172 9.376 8.137
591 11.987 10.246 11.428 10.151
596 8.990 8.133 9.147 8.059
All 10.324 8.873 10.425 8.758
The NHITS-PK model effectively learns the inhibitory treatment effects of insulin on blood glucose.
We examine whether the NHITS-PK model can accurately capture treatment effects of insulin on blood glucose by generating forecasts for three scenarios: the original bolus insulin doses, bolus insulin doses removed, and bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 8. This result underscores the importance of models that can effectively capture treatment effects, enabling them to adapt to new treatment doses for more reliable and accurate forecasts.
Figure 7: Blood glucose forecasts of the final horizon point during 5 randomly selected hyperglycemic periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients. Forecasts for the NHITS-PK model (orange, dashed) are more aligned with ground truth (black, solid) compared to the NHITS sparse exogenous model blue, dash-dotted).
Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed). Each point represents the final forecasted timestamp from a rolling window forecast.
@@ 6 Discussion
The proposed PK encoder can substantially reduce forecasting errors in deep learning models for both simulated and real-world data. These improvements are not only statistically significant, but they have potential for positive real-world clinical impact. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts these events earlier, up to 3.6 minutes compared to the NHITS sparse exogenous baseline model, for individual OhioT1DM patients, as shown in Fig. 9. This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations. Improving the prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009). Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing. While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge. It can also be easily adapted for other types of medication administration, such as oral or intravenous injections.
In addition to improving performance over global model counterparts, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models. This result is of substantial importance in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts. Moreover, our proposed hybrid global-local model offers practical benefits beyond improved forecasting performance as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took approximately seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. The hybrid model also has the potential to be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
@@
Our study demonstrates that leveraging domain knowledge to integrate exogenous variables in models offers performance benefits over using sparse features or the “Sum Total” feature engineering approach, both of which are outperformed by our proposed deep PK encoding method. Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
Limitations and Future Work.
This study has two limitations related to the PK encoder: we assume that the concentration curves exhibit 100% bioavailability and maintain a consistent peak concentration across patients. These simplifying assumptions are intended to focus model learning on the parameter $k$, rather than on the area or peak of the concentration curve, to emphasize the spread or duration of the concentration curve and to reduce potential issues with parameter identification. Future work will explore expanding concentration curve parameters to estimate area and peak parameters to potentially enable more flexible concentration curves. We will also aim to apply our approach to additional cohorts of patients, such as those with T2DM, as more data become available, to further validate our findings. Future work will also explore modeling dose-dependent effects using more dynamic nonlinear PK frameworks.
Figure 9: The PK encoder enables improved NHITS model prediction lead times for hyperglycemic events (CGM $\geq$ 180) for 9 of 12 OhioT1DM patients.
@@
Data and Code Availability
We use open-source simulated and real-world data that is available to other researchers. The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository (Xie, 2018; Visentin et al., 2016). The real-world data includes the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu (2020). Implementations of models used in our work were obtained from the open-source Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries. We implemented our method by adopting the Neuralforecast repository to leverage their model training framework. Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.